OpenStudy (anonymous):

looking for the discriminant (b^ - 4ac), if the formula is 4x^ + 4x + 1 = 0 am I right in thinking this will become, 4x^ - 4 x 16 x 1 which equals -48 and therefore there are no solutions?

OpenStudy (turingtest):

not quite... what is b ? what is a ? what is c ?

OpenStudy (anonymous):

a = 4x squared b = 4x c= 1

OpenStudy (turingtest):

no, a, b, and c are just the coefficients (the numbers) do not include the x part

OpenStudy (anonymous):

a = 4 b = 4 c = 1

OpenStudy (turingtest):

so now what do you get for the discriminant?

OpenStudy (anonymous):

- 48?

OpenStudy (turingtest):

4^2-4(4)(1)=-48 ? I don't think so, recheck your arithmetic

OpenStudy (anonymous):

I'm happy with the first part, 4^2 = 16 I think B is confusing me is it -4x4 which is -16 + 1(c) so total answer = 1?

OpenStudy (turingtest):

4^2=16 4(4)(1)=16 so 4^2-4(4)(1)=16-16

OpenStudy (anonymous):

of course because your multiplying AxBxC so 16- 16x1 = 0 so back to the original question now you've helped me with this as the discriminant is now 0 that means there can only be one solution? correct?

OpenStudy (turingtest):

yes, a "double root" of one real number

OpenStudy (anonymous):

thank you for your assistance Master Maths Jedi! :)

OpenStudy (turingtest):

ah, very welcome you are force be with you ;)